# On the asymptotic profile of solutions to semilinear damped wave equations with critical nonlinearities

### v1

@@ Abstract.
Dinh Van Duong1,∗, Trung Loc Tang2
1Faculty of Mathematics and Informatics, Hanoi University of Science and Technology
No.1 Dai Co Viet road, Hanoi, Vietnam
@@ 1. Introduction
In this work, we consider the following Cauchy problem for the semilinear damped wave equation:
@@
In addition, they concluded that the damped wave equation has a diffusive structure as $t\to\infty$. This phenomenon has been further investigated in many subsequent papers, such as [6, 23, 31, 14, 2] and the references therein. From this, they conclude that the diffusion phenomenon bridges decay properties of solutions to the Cauchy problem for the classical damped wave equation (3) and solutions to the following Cauchy problem for the heat equation:
@@
Specifically, they obtain the approximation of the solution $u$ by the Gauss kernel when $t$ is large
@@
More extended, Michihisa [23] discovered the higher-order asymptotic behavior of solution to problem (3) in the $L^{2}$ framework for any spatial dimension $n$ by using some tools including Taylor series expansion and Faà di Bruno’s formula. Takeda [31] found the higher-order asymptotic behavior of solution to problem (3) in $L^{q}$ framework for some $q\in(1,\infty)$ and concluded that the effect of $u_{tt}$ in (3) is not negligible for the third-order expansion.
For the semilinear problem (2) under the additional regularity $L^{1}$ for the initial data, in [15, 32, 16, 18], the authors showed that the Fujita exponent $p_{F}(n):=1+2/n$ is the critical exponent of (2) and the paper [33] showed that the case $p=p_{F}(n)$ belongs to the blow-up range. Here, the critical exponent is understood as the threshold between the global (in time) existence of small data mild solutions and the blow-up of solutions even for small data. The Fujita exponent $p_{F}$ is also the critical exponent for the semilinear problem (4), whose right-hand side is the nonlinear term $|v|^{p}$. Specifically, Ikehata et al. [15] succeeded in proving the global existence and optimal decay estimates of the total energy of the weak solutions to problem (2) with the power $p>p_{F}(n)$ for $n=1,2$, and $p>2$ for $n\geq 3$. It should be mentioned that the results in [32] fully depend on the compactness assumption on the support of initial data, while in [15] these were removed. Next, Gallay-Raugel [10] proved that global solutions of nonlinear damped wave equation behaves like those of nonlinear heat equations with suitable data, including more general nonlinearity for $n=1$. Karch [21] proved the approximation of the solution to (2) by the Gauss kernel for $p\geq 1+4/n$. Nishihara [27] proved it for $p>1+2/n$ when $n=3$, followed by [26] for $n=1$, [20] for $n=2$, [28] for $n=4,5$ and [19] for all $n\geq 1$. More gererally, Kawakami-Takeda [22] established the higher order asymptotic expansion of the solution to (2) under suitable assumptions for the nonlinearity and the initial data. For the blow-up range $p\leq 1+2/n$, according to the works [24, 13, 25], the sharp lifespan estimates for blow-up solutions to (2) in all spatial dimensions have been investigated. Here, we denote that $T_{\varepsilon}$ is the lifespan of the local solution in the following sense:
@@
Consequently, these papers provided the following sharp lifespan estimate for blow-up solutions:
$T(\varepsilon)\sim\begin{cases}\vskip 5.69046ptC\varepsilon^{-\frac{2(p-1)}{2-n(p-1)}}&\text{ if }1<p<1+\displaystyle\frac{2}{n},\\ \exp\big(C\varepsilon^{-\frac{2}{n}}\big)&\text{ if }p=1+\displaystyle\frac{2}{n},\end{cases}$
where $C$ is a positive constant independent of $\varepsilon$.
Under additional regularity $L^{m}$ for the initial data, with $m\in[1,2]$, papers [16, 7] and references therein identified the critical exponent of problem (2) as $p_{c}(m)=p_{F}(n/m)=1+2m/n.$ However, the authors did not provide conclusions regarding the solution’s properties when $p=p_{c}(m)$. Paper [14] has proved the existence of global mild solutions when $p=p_{c}(m)$ for all $m\in(1,2]$ in small dimentional spaces along with some other conditions. In addition, they proved that the global solution is approximated by that of the linear heat equation (4) for all dimension spaces $n\geq 1$.
From the above discussion, we can see that the Fujita exponent $p_{F}=1+2/n$ serves as the threshold separating the blow-up range from the global existence range of global (in time) solutions to problem (2) with small data. Moreover, the critical case $p=p_{F}$ belongs to the blow-up range. Recently, Ebert et al. [8] investigated the Cauchy problem (1) with nonlinear term $|u|^{1+\frac{2}{n}}\mu(|u|)$ and established the sharp condition on $\mu(s)$ that separates the blow-up case from the global existence case. Specifically, the problem (1) has a unique global (in time) small data solution if $n=1,2$ and the Dini condition satisfied. Moreover, problem (1) admits a blow-up solution in finite time for all spatial dimensions $n\geq 1$ if $\mu$ satisfies the non-Dini condition.
@@
Extending the work [8], Girardi [11] showed that the results obtained in [8] can be generalized to more general semilinear evolution models with the Fujita-type critical exponent; in particular, the critical behavior of nonlinearity is still described by the Dini condition (5). From this, they extended the global existence result of (1) to the case $n=3$ by using the estimates of Nishihara in [27]. Continuing with this topic, Dao-Reissig [4] studied for semilinear damped wave systems, while the papers of Chen-Girardi [3] and Dao-Son [5] investigated it in the framework of evolution equations. They also proved that, under the condition (5), these problems admit a unique global (in time) solution for small data. However, their assumptions on the spatial dimension are rather restrictive, since they are constrained by tools from harmonic analysis. Therefore, our first main goal in this paper not only extends the global existence property of (1) to dimensions $1\leq n\leq 4$ under condition (5) by applying the linear estimates obtained in [14] , but also provides a new approach to relax the dimensional restrictions in [4, 5, 3]. Next, our second main objective is to determine the asymptotic behavior of global (in time) solutions to (1) in spatial dimensions $1\leq n\leq 4$. More specifically, we show that the behavior of the solution at $t\to\infty$ is identified by the Gauss kernel under the Dini condition (5) for the global existence
@@
for all $q\in\left[\min\{2,1+2/n\},+\infty\right]$ and $M$ defined by (8). Finally, we will provide sharp lifespan estimates for local solutions of problem (1) under the non-Dini condition (6) in spatial dimensions $1\leq n\leq 4$.
@@
  • As usual, $H^{a}_{m}$ and $\dot{H}^{a}_{m}$, with $m\in(1,\infty),a\geq 0$, denote potential spaces based on $L^{m}$ spaces. Here $\big<\nabla\big>^{a}$ and $|\nabla|^{a}$ stand for the pseudo-differential operators with symbols $\big<\xi\big>^{a}$ and $|\xi|^{a}$, respectively, where the symbol $\langle x\rangle:=\sqrt{|x|^{2}+1}$ denotes the Japanese bracket.
  • For later convenience, we denote by $C$ and $C_{\ell}$ with $\ell\in\mathbb{N}$ positive constants independent of $t$, which may be changed from line to line.
Main results. Let us state the global (in time) existence of small data solutions, along with asymptotic behaviors for global solution and sharp lifespan estimates for blow-up solutions , which will be proved in this paper.
@@
Consider that the assumptions of Theorem 1.1 hold, the global (in time) small data solution to (1) satisfies the following estimates for $t\gg 1$:
@@
From the statements of Theorem 1.2, we obtain the asymptotic behavior of the global solution at $t\to\infty$ is identified by the Gauss kernel under the Dini condition (5) as follow:
@@
where $t\gg 1$. These estimates can be easily derived from the conclusion of Theorem 1.2 together with the triangle inequality.
If the non-Dini condition (6) occurs, then Theorem 5 in [8] shows that the solution to (1) blows up in finite time. Now, we define the following function:
Theorem 1.3 (Sharp lifespan estimates for blow-up solutions).
[…]
$\int_{\mathbb{R}^{n}}(u_{0}(x)+u_{1}(x))dx>0.$ (11)
$T_{\varepsilon}\lesssim\Psi^{-1}\left(C\varepsilon^{-\frac{2}{n}}\right)$
satisfies the lower bound estimate
$T_{\varepsilon}\gtrsim\Psi^{-1}\left(c\varepsilon^{-\frac{2}{n}}\right),$
Remark 1.4.
Linking the achieved estimates in Theorem 1.3, one recognizes that the sharp estimates for the lifespan $T_{\varepsilon}$ of blow-up solutions to the Cauchy problem (1) in spatial dimensions $1\leq n\leq 4$ under the non-Dini condition are determined by the following relations:
$T_{\varepsilon}\sim\Psi^{-1}\big(c\varepsilon^{-\frac{2}{n}}).$
This is also a new result for spatial dimensional spaces $1\leq n\leq 4$.
This paper is organized as follows: In Section 2, we provide the proof of global (in time) existence results for solutions to the problem (1). Subsequently, in Section 3, we describe the asymptotic behavior of the global solution to (1). Finally, we establish sharp lifespan estimates for blow-up solutions in Section 4.
@@ 2. Global Existence
2.1. Philosophy of our approach
@@
Now, we restate the important result as follows.
@@
Due to the similarity in the proof and the fact that the result is completely analogous to Theorem 1.1 in [14], we present only the proof of the case $j=0$. Specifically, using Proposition 2.4 in [14], we gain
$\||\nabla|^{s_{1}}\chi_{\rm L}(|\nabla|)\mathcal{K}(t,x)\ast_{x}\varphi(x)\|_{L^{q}}\lesssim(1+t)^{-\frac{n}{2}(\frac{1}{\rho}-\frac{1}{q})-\frac{s_{1}-s_{2}}{2}}\||\nabla|^{s_{2}}\chi_{\rm L}(|\nabla|)\varphi\|_{L^{\rho}},\text{ for all }1\leq\rho\leq q\leq\infty.$
@@
$\quad\lesssim\||\nabla|^{s_{1}}\langle\nabla\rangle^{d}\chi_{\rm H}(|\nabla|)\mathcal{K}(t,x)\ast_{x}\varphi(x)\|_{L^{r}}\lesssim e^{-ct}\||\nabla|^{s_{1}+d}\chi_{\rm H}(|\nabla|)\varphi\|_{H^{\beta_{r}-1}_{r}},$
for all $1<r<\infty$ and $d>n/r$. This completes the proof. ∎
@@
$\|\varphi\|_{X(T)}:=$ $\sup_{t\in[0,T]}\bigg\{(1+t)^{\frac{n}{2}(1-\frac{1}{\alpha})}\|\varphi(t,\cdot)\|_{L^{\alpha}}+(1+t)^{\frac{n}{4}+1}\|\varphi(t,\cdot)\|_{\dot{H}^{2}}+(1+t)^{\frac{n}{2}}\|\varphi(t,\cdot)\|_{L^{\infty}}\bigg\}$
@@
Firstly, we can easily see that $X(T,M)\subset Y(T)$ by interpolation. Therefore, it suffices to show that if a sequence in $X(T,M)$ converging in $Y(T)$, its limit will belong to $X(T,M)$. More specifically, we assume that $\{\varphi_{j}\}_{j=1}^{\infty}\subset X(T,M)$ and $\varphi_{j}\to\varphi\in Y(T)$ as $j\to\infty$. It is a fact that
@@
Next, we define the operator $\Phi$ on the space $X(T)$
@@
Let $1<p,\,p_{0},\,p_{1}<\infty$, $a>0$ and $\theta\in[0,a)$. Then, it holds the following fractional Gagliardo-Nirenberg inequality:
$\|u\|_{\dot{H}^{\theta}_{p}}\lesssim\|u\|_{L^{p_{0}}}^{1-\omega(\theta,a)}\,\|u\|_{\dot{H}^{a}_{p_{1}}}^{\omega(\theta,a)},$
where $\omega(\theta,a)=\displaystyle\frac{\frac{1}{p_{0}}-\frac{1}{p}+\frac{\theta}{n}}{\frac{1}{p_{0}}-\frac{1}{p_{1}}+\frac{a}{n}}$ and $\frac{\theta}{a}\leq\omega(\theta,a)\leq 1$.
@@
Firstly, we will prove the following auxiliary results.
@@
$\|\nabla\mathcal{N}(u(\tau,\cdot))\|_{L^{\theta}}$ $\lesssim\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)(1+\tau)^{-\frac{n}{2}(1+\frac{2}{n}-\frac{1}{\theta})-\frac{1}{2}}\|u\|_{X(T)}^{1+\frac{2}{n}}\quad\text{ for }\theta\in\left[\alpha,\frac{2n}{[n-2]^{+}}\right).$
@@
Firstly, applying interpolation, we can easily see that
@@
Therefore, we can obtain an immediate result
@@
$\omega_{1}:=\displaystyle\frac{\displaystyle\frac{1}{\alpha}-\frac{1}{\theta}+\frac{1}{n}}{\displaystyle\frac{1}{\alpha}-\frac{1}{2}+\frac{2}{n}}\in[0,1]\quad\text{ for all }\quad\theta\in\left[\alpha,\frac{2n}{[n-2]^{+}}\right).$
Therefore, we get
@@
where we note that $1+\tau\geq 1+t-\tau$ for all $\tau\in[t/2,t]$. By using the change of variables $\theta_{1}=\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}$ and $\theta_{2}=\varepsilon_{0}(1+t-\tau)^{-\frac{n}{2}}$, we can easily see that
$\int_{0}^{t/2}(1+\tau)^{-1}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau+\int_{t/2}^{t}(1+t-\tau)^{-1}\mu\left(\varepsilon_{0}(1+t-\tau)^{-\frac{n}{2}}\right)d\tau\lesssim\mathcal{I}(\varepsilon_{0}).$
@@
where we choose $d:=n/r+\delta$ with $\delta$ is a sufficiently small constant. Noting that the conditions $r\in\left(2,\displaystyle\frac{2n}{[n-2]^{+}}\right)$ and $1\leq n\leq 4$ imply
$0<d+\beta_{r}=\frac{n}{r}+\delta+(n-1)\left(\frac{1}{2}-\frac{1}{r}\right)<2,$
@@
$\|\mathcal{N}(u(\tau,\cdot))\|_{H_{r}^{d+\beta_{r}-1}}$ $\lesssim\|\mathcal{N}(u(\tau,\cdot))\|_{H_{r}^{1}}\approx\|\mathcal{N}(u(\tau,\cdot))\|_{L^{r}}+\|\nabla\mathcal{N}(u(\tau,\cdot))\|_{L^{r}}$
@@
due to using again Lemma 2.3. As a result, we get
@@
$\quad+(1+t)^{-\frac{n}{2}}\|u\|_{X(T)}^{1+\frac{2}{n}}\int_{t/2}^{t}(1+t-\tau)^{-1}\mu\left(\varepsilon_{0}(1+t-\tau)^{-\frac{n}{2}}\right)d\tau$
$\lesssim\mathcal{I}(\varepsilon_{0})(1+t)^{-\frac{n}{2}}\|u\|_{X(T)}^{1+\frac{2}{n}}.$
From this, we obtain the following estimate:
$\|u(t,\cdot)\|_{L^{\infty}}\lesssim\mathcal{I}(\varepsilon_{0})(1+t)^{-\frac{n}{2}}\|u\|_{X(T)}^{1+\frac{2}{n}}.$ (14)
Finally, we need to estimate $\|u^{\rm non}(\tau,\cdot)\|_{L^{\alpha}}$. We note that $\beta_{\alpha}<1$ for all $1\leq n\leq 4$. Therefore, employing again Lemma 2.3 combined with Lemma 2.1 with $j=0,\,\rho=1,q=\alpha,(s_{1},s_{2})=(0,0)$ for $\tau\in[0,t/2)$ and $\rho=q=\alpha,(s_{1},s_{2})=(0,0)$ for $\tau\in[t/2,t]$ we have
@@
$\lesssim(1+t)^{-\frac{n}{2}(1-\frac{1}{\alpha})}\|u\|_{X(T)}^{1+\frac{2}{n}}\int_{0}^{t/2}(1+\tau)^{-1}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau$
$\quad+(1+t)^{-\frac{n}{2}(1-\frac{1}{\alpha})}\|u\|_{X(T)}^{1+\frac{2}{n}}\int_{t/2}^{t}(1+t-\tau)^{-1}\mu\left(\varepsilon_{0}(1+t-\tau)^{-\frac{n}{2}}\right)d\tau$
$\lesssim\mathcal{I}(\varepsilon_{0})(1+t)^{-\frac{n}{2}(1-\frac{1}{\alpha})}\|u\|_{X(T)}^{1+\frac{2}{n}}.$
In summary, we have the following estimate:
$\|u^{\rm non}(t,\cdot)\|_{L^{\alpha}}\lesssim\mathcal{I}(\varepsilon_{0})(1+t)^{-\frac{n}{2}(1-\frac{1}{\alpha})}\|u\|_{X(T)}^{1+\frac{2}{n}}.$ (15)
@@
$\|\Phi[u]-\Phi[v]\|_{Y(T)}$ $\lesssim\sup_{t\in[0,T]}\left\{\int_{0}^{t}(1+t-\tau)^{-\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\|\mathcal{N}(u(\tau,\cdot))-\mathcal{N}(v(\tau,\cdot))\|_{L^{m}\cap H^{-1}}d\tau\right\}$
$\lesssim\sup_{t\in[0,T]}\left\{\int_{0}^{t}(1+t-\tau)^{-\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\|\mathcal{N}(u(\tau,\cdot))-\mathcal{N}(v(\tau,\cdot))\|_{L^{m}}d\tau\right\},$

### v2

@@ Abstract.
Trung Loc Tang1, Dinh Van Duong2,∗
@@ 1 Department of Mathematics, Hanoi National University of Education
2Faculty of Mathematics and Informatics, Hanoi University of Science and Technology
No.1 Dai Co Viet road, Hanoi, Vietnam
@@ 1. Introduction
In this work, we consider the following Cauchy problem for semilinear damped wave equations:
@@
In addition, they concluded that the damped wave equation has a diffusive structure as $t\to\infty$. This phenomenon has been further investigated in many subsequent papers, such as [7, 24, 32, 15, 2] and the references therein. From this, they conclude that the diffusion phenomenon bridges the decay properties of solutions to the Cauchy problem for the classical damped wave equation (3) and solutions to the following Cauchy problem for the heat equation:
@@
Specifically, they obtain the following approximation of the solution $u$ by the Gauss kernel when $t$ is large:
@@
Furthermore, Michihisa [24] discovered the higher-order asymptotic behavior of the solution to (3) in the $L^{2}$ framework for any spatial dimension $n$ by using some tools, including Taylor series expansion and Faà di Bruno’s formula. Takeda [32] found the higher-order asymptotic behavior of solution to (3) in the $L^{q}$ framework for some $q\in(1,\infty)$ and concluded that the effect of $u_{tt}$ in (3) is not negligible for the third-order expansion.
Taking into account the semilinear problem (2) under the additional regularity $L^{1}$ for the initial data, in [16, 33, 17, 19], the authors showed that the Fujita exponent $p_{F}(n):=1+2/n$ is the critical exponent of (2) and the paper [34] showed that the value $p=p_{F}(n)$ belongs to the blow-up range. Here, the critical exponent is understood as the threshold between the global (in time) existence of small data solutions and the blow-up of solutions even for small data. The Fujita exponent $p_{F}$ is also the critical exponent for the semilinear problem (4), whose right-hand side is the nonlinear term $|v|^{p}$. Specifically, Ikehata et al. [16] succeeded in proving the global existence and optimal decay estimates of the total energy of the weak solutions to problem (2) with the power $p>p_{F}(n)$ for $n=1,2$, and $p>2$ for $n\geq 3$. It should be mentioned that the results in [33] fully depend on the compactness assumption on the support of initial data, while in [16] these were removed. Next, Gallay-Raugel [11] proved that global solutions of nonlinear damped wave equation behaves like those of nonlinear heat equations with suitable data, including more general nonlinearity for $n=1$. Karch [22] proved the approximation of the solution to (2) by the Gauss kernel for $p\geq 1+4/n$. Nishihara [28] proved it for $p>1+2/n$ when $n=3$, followed by [27] for $n=1$, [21] for $n=2$, [29] for $n=4,5$ and [20] for all $n\geq 1$. More gererally, Kawakami-Takeda [23] established the higher order asymptotic expansion of the solution to (2) under suitable assumptions for the nonlinearity and the initial data. For the blow-up range $p\leq 1+2/n$, according to the works [25, 14, 26], the sharp lifespan estimates for blow-up solutions to (2) in all spatial dimensions have been investigated. Here, we denote by $T_{\varepsilon}$ the lifespan of solution in the following sense:
@@
Consequently, these papers provided the following sharp lifespan estimates:
$T_{\varepsilon}\sim\varepsilon^{-\frac{2(p-1)}{2-n(p-1)}}\text{ if }1<p<1+\displaystyle\frac{2}{n}\quad\text{ and }\quad\log(T_{\varepsilon})\sim\varepsilon^{-\frac{2}{n}}\text{ if }p=1+\displaystyle\frac{2}{n}.$
From the above discussion, one has the Fujita exponent $p_{F}=1+2/n$ serves as the threshold separating the blow-up range from the global existence range of global (in time) solutions to problem (2) with small data. Moreover, the critical case $p=p_{F}$ belongs to the blow-up range. Recently, Ebert et al. [9] investigated the Cauchy problem (1) with nonlinear term $|u|^{1+\frac{2}{n}}\mu(|u|)$ and established the sharp condition on $\mu(s)$ that separates the blow-up case from the global existence case. Specifically, the problem (1) has a unique global (in time) small data solution if $n=1,2$ and the Dini condition is satisfied. Moreover, problem (1) admits a blow-up solution in finite time for all spatial dimensions $n\geq 1$ if $\mu$ satisfies the non-Dini condition.
@@
Extending the work [9], Girardi [12] showed that the results obtained in [9] can be generalized to more general semilinear evolution models with the Fujita-type critical exponent; in particular, the critical behavior of nonlinearity is still described by the Dini condition (5). From this, they extended the global existence result of (1) to the case $n=3$ by using the estimates of Nishihara in [28]. Continuing with this topic, Dao-Reissig [5] studied for semilinear damped wave systems, while the papers of Chen-Girardi [4] and Dao-Son [6] investigated it in the framework of evolution equations. They also proved that, under the condition (5), these problems admit a unique global (in time) solution for small data. However, their assumptions on the spatial dimension are rather restrictive, since they are constrained by tools from harmonic analysis. In addition, the readers may refer to paper [3] for the semilinear classical wave equations. Our first main goal in this paper extends the global existence property of (1) to dimensions $1\leq n\leq 4$ under condition (5) by applying the linear estimates obtained in [15]. Next, our second main objective is to determine the asymptotic behavior of global (in time) solutions to (1) in spatial dimensions $1\leq n\leq 4$. More specifically, we show that the behavior of the solution at $t\to\infty$ is identified by the Gauss kernel under the Dini condition (5) for the global existence
@@
for all $q\in\left[\min\{2,1+2/n\},+\infty\right]$ and $M$ defined by (8). Finally, we will provide some remarks on sharp lifespan estimates for blow-up solutions to problem (1) under the non-Dini condition (6).
@@
  • As usual, $H^{a}_{m}$ and $\dot{H}^{a}_{m}$, with $m\in(1,\infty),a\geq 0$, denote potential spaces based on $L^{m}$ spaces. Here $\langle\nabla\rangle^{a}$ and $|\nabla|^{a}$ stand for the pseudo-differential operators with symbols $\big<\xi\big>^{a}$ and $|\xi|^{a}$, respectively, where the symbol $\langle x\rangle:=\sqrt{|x|^{2}+1}$ denotes the Japanese bracket.
Main results. Let us state the global (in time) existence of small data solutions, together with their asymptotic behavior, which will be proved in this paper.
@@
Assume that the assumptions of Theorem 1.1 hold, the global (in time) small data solution to (1) satisfies the following estimates for $t\gg 1$:
@@
From the statements of Theorem 1.2, we see that the asymptotic behavior of the global solution at $t\to\infty$ is identified by the Gauss kernel under the Dini condition (5) as follows:
@@
where $t\gg 1$. These estimates can be derived from the conclusion of Theorem 1.2 together with the triangle inequality.
This paper is organized as follows: In Section 2, we provide the proof of global (in time) existence results for solutions to the problem (1). Subsequently, in Section 3, we describe the asymptotic behavior of the global solution to (1). Finally, we provide some remarks on sharp lifespan estimates in Section 4.
@@ 2. Global Existence
2.1. Linear estimates and solution spaces
@@
Now, we restate the following important result.
@@
Due to the similarity in the proof, we present only the proof of the case $j=0$. Specifically, using Proposition 2.4 in [15], we gain
$\||\nabla|^{s_{1}}\chi_{\rm L}(|\nabla|)\mathcal{K}(t,x)\ast_{x}\varphi(x)\|_{L^{q}}$
$\quad\lesssim(1+t)^{-\frac{n}{2}(\frac{1}{\rho}-\frac{1}{q})-\frac{s_{1}-s_{2}}{2}}\||\nabla|^{s_{2}}\chi_{\rm L}(|\nabla|)\varphi\|_{L^{\rho}},\text{ for all }1\leq\rho\leq q\leq\infty.$
@@
$\quad\lesssim\||\nabla|^{s_{1}}\langle\nabla\rangle^{d}\chi_{\rm H}(|\nabla|)\mathcal{K}(t,x)\ast_{x}\varphi(x)\|_{L^{r}}\lesssim e^{-ct}\||\nabla|^{s_{1}+d}\chi_{\rm H}(|\nabla|)\varphi\|_{H^{\beta_{r}-1}_{r}},$ (10)
for all $1<r<\infty$ and $d>n/r$. These complete the proof of case $j=0$. ∎
@@
$\|\varphi\|_{X(T)}$
$\,\,:=\sup_{t\in[0,T]}\bigg\{(1+t)^{\frac{n}{2}(1-\frac{1}{\alpha})}\|\varphi(t,\cdot)\|_{L^{\alpha}}+(1+t)^{\frac{n}{4}+1}\|\varphi(t,\cdot)\|_{\dot{H}^{2}}+(1+t)^{\frac{n}{2}}\|\varphi(t,\cdot)\|_{L^{\infty}}\bigg\}$
@@
Firstly, we can see that $X(T,M)\subset Y(T)$ by interpolation. Therefore, it suffices to show that if a sequence in $X(T,M)$ converging in $Y(T)$, its limit will belong to $X(T,M)$. More specifically, we assume that $\{\varphi_{j}\}_{j=1}^{\infty}\subset X(T,M)$ and $\varphi_{j}\to\varphi\in Y(T)$ as $j\to\infty$. One has the following relation:
@@
Next, we consider the operator $\Phi$ on the space $X(T)$
@@
Let $1<p,\,p_{0},\,p_{1}<\infty$, $a>0$ and $\sigma\in[0,a)$. Then, we have the following fractional Gagliardo-Nirenberg inequality:
$\|u\|_{\dot{H}^{\sigma}_{p}}\lesssim\|u\|_{L^{p_{0}}}^{1-\omega(\sigma,a)}\,\|u\|_{\dot{H}^{a}_{p_{1}}}^{\omega(\sigma,a)},$
where $\omega(\sigma,a):=\displaystyle\frac{\frac{1}{p_{0}}-\frac{1}{p}+\frac{\sigma}{n}}{\frac{1}{p_{0}}-\frac{1}{p_{1}}+\frac{a}{n}}$ and $\frac{\sigma}{a}\leq\omega(\sigma,a)\leq 1$.
@@
Firstly, we will prove the following results.
@@
$\|\nabla\mathcal{N}(u(\tau,\cdot))\|_{L^{\theta}}$ $\lesssim\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)(1+\tau)^{-\frac{n}{2}(1+\frac{2}{n}-\frac{1}{\theta})-\frac{1}{2}}\|u\|_{X(T)}^{1+\frac{2}{n}}\quad\text{ for }\theta\in\left[2,\frac{2n}{[n-2]^{+}}\right).$
@@
Firstly, applying interpolation to have
@@
Therefore, we obtain the following relation:
@@
$\omega_{1}:=\displaystyle\frac{\displaystyle\frac{1}{\alpha}-\frac{1}{\theta}+\frac{1}{n}}{\displaystyle\frac{1}{\alpha}-\frac{1}{2}+\frac{2}{n}}\in\left[\frac{1}{2},1\right]\quad\text{ for all }\quad\theta\in\left[2,\frac{2n}{[n-2]^{+}}\right).$
From this, we get
@@
$\lesssim\|u\|_{X(T)}^{1+\frac{2}{n}}\int_{0}^{t/2}(1+t-\tau)^{-\frac{n}{4}-1}(1+\tau)^{-1}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau$
$\qquad+\|u\|_{X(T)}^{1+\frac{2}{n}}\int_{t/2}^{t}(1+t-\tau)^{-\frac{1}{2}}(1+\tau)^{-\frac{n}{4}-\frac{3}{2}}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau$
@@
where noting that $1+t\sim 1+\tau\geq 1+t-\tau$ for all $\tau\in[t/2,t]$. Using the change of variables $\theta_{1}=\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}$ and $\theta_{2}=\varepsilon_{0}(1+t-\tau)^{-\frac{n}{2}}$, we get
$\int_{0}^{t/2}(1+\tau)^{-1}\mu\left(\varepsilon_{0}(1+\tau)^{-\frac{n}{2}}\right)d\tau+\int_{t/2}^{t}(1+t-\tau)^{-1}\mu\left(\varepsilon_{0}(1+t-\tau)^{-\frac{n}{2}}\right)d\tau\lesssim\mathcal{I}(\varepsilon_{0}).$ (12)
@@
where $d:=n/r+\delta$ with $\delta$ is a sufficiently small constant. Noting that the conditions $r\in\left(2,\displaystyle\frac{2n}{[n-2]^{+}}\right)$ and $1\leq n\leq 4$ imply
$0<d+\beta_{r}=\frac{n}{r}+\delta+(n-1)\left(\frac{1}{2}-\frac{1}{r}\right)<2,$ (14)
@@
$\|\mathcal{N}(u(\tau,\cdot))\|_{H_{r}^{d+\beta_{r}-1}}$ $\lesssim\|\mathcal{N}(u(\tau,\cdot))\|_{H_{r}^{1}}\sim\|\mathcal{N}(u(\tau,\cdot))\|_{L^{r}}+\|\nabla\mathcal{N}(u(\tau,\cdot))\|_{L^{r}}$
@@
due to using again Lemma 2.3. Combining this with the relation (12) and $n<2r$, we get
@@
$\lesssim\mathcal{I}(\varepsilon_{0})(1+t)^{-\frac{n}{2}}\|u\|_{X(T)}^{1+\frac{2}{n}}.$ (15)
Finally, we need to estimate $\|u^{\rm non}(\tau,\cdot)\|_{L^{\alpha}}$. We have $\beta_{\alpha}<1$ for all $1\leq n\leq 4$. Therefore, employing again Lemma 2.3 combined with Lemma 2.1 with $j=0,\,\rho=1,q=\alpha,(s_{1},s_{2})=(0,0)$ for $\tau\in[0,t/2)$ and $\rho=q=\alpha,(s_{1},s_{2})=(0,0)$ for $\tau\in[t/2,t]$ we have
@@
$\lesssim\mathcal{I}(\varepsilon_{0})(1+t)^{-\frac{n}{2}(1-\frac{1}{\alpha})}\|u\|_{X(T)}^{1+\frac{2}{n}}.$ (16)
@@
$\|\Phi[u]-\Phi[v]\|_{Y(T)}$
$\quad\lesssim\sup_{t\in[0,T]}\left\{\int_{0}^{t}(1+t-\tau)^{-\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\|\mathcal{N}(u(\tau,\cdot))-\mathcal{N}(v(\tau,\cdot))\|_{L^{m}\cap H^{-1}}d\tau\right\}$
$\quad\lesssim\sup_{t\in[0,T]}\left\{\int_{0}^{t}(1+t-\tau)^{-\frac{n}{2}(\frac{1}{m}-\frac{1}{2})}\|\mathcal{N}(u(\tau,\cdot))-\mathcal{N}(v(\tau,\cdot))\|_{L^{m}}d\tau\right\},$